(10 pts) A firm has the following relationship between output (Q) and total cost (TC): Q...

  1. (10 pts) A firm has the following relationship between output (Q) and total cost (TC):

Q

TC

0

$100

1

110

2

130

3

160

4

200

5

250

6

310

7

380

8

460

9

550

10

650

  1. Say the firm is a perfect competitor. If the market price for its product is $ 80, at what output level will this firm produce at (as a profit maximizer)?
  1. At the output level in (a), are firms in this industry making a profit or loss? Will firms enter or exit the market?
  1. Say this firm is a monopoly. If the firm maximizes profit where marginal revenue equals $ 50, at what output level will the firm be producing at (as a profit maximizer)?
Answer #1

MC = Change in TC / Change in Q

Output TC MC
0 100
1 110 10
2 130 20
3 160 30
4 200 40
5 250 50
6 310 60
7 380 70
8 460 80
9 550 90
10 650 100

a. In perfect competition profit is maximized when P = MC

Here P = MC = 80 at output level of 8

b. Profit / Loss = TR - TC = 8 * 80 - 460 = 640 - 460 = 180

As Firm is earning positive profit, firms will enter the market

c. In a monopoly, profit is maximized where MR = MC

Here MR = MC = 50 at output level of 5